1. Find the time required for an investment of 5000 dollars to grow to 7500 dollars at an interest rate of 7.5 percent per year, compounded quarterly.
Round your answer to two decimal places
Your answer is t=

2. You have $5,000 on a credit card that charges a 23% interest rate. If you want to pay off the credit card in 4 years, how much will you need to pay each month (assuming you don't charge anything new to the card)?

3. Suppose you want to have $300,000 for retirement in 20 years. Your account earns 7% interest. How much would you need to deposit in the account each month?

$_______?

$ ______each month
Answer #1

1. Information provided:

Present value= $5,000

Future value= $7,500

Interest rate= 7.50%/4= 1.8750% per quarter

The time is computed by entering the below in a financial calculator:

PV= -5,000

FV= 7,500

I/Y= 1.8750

Press the CPT key and N to compute the time.

The value obtained is 21.8269.

Therefore, $5,000 will take 21.83 quarters to grow t0 $7,500.

In case of any query, kindly comment on the solution.

2.Information provided:

Present value= $5,000

Interest rate= 23%/12= 1.9167% per month

Time= 4 years*12= 48 months

The monthly payment to pay off the credit card is calculated by entering the below in a financial calculator:

PV= -5,000

I/Y= 1.9167

N= 48

Press the CPT key and PMT to compute the monthly payment.

The value obtained is 160.2574.

Therefore, I would need to pay $160.26 each month.

3.Information provided:

Future value= $300,000

Time= 20 years*12= 240 months

Interest rate= 7%/12= 0.5833% per month

The amount to be deposited each month is is calculated by entering the below in a financial calculator:

FV= 300,000

N= 240

I/Y= 0.5833

The monthly deposit is calculated by entering the below in a financial calculator:

FV= 300,000

N= 240

I/Y= 0.5833

Press the CPT key and PMT to compute the monthly payment.

The value obtained is 575.8968.

Therefore, I would need to deposit $575.90 each month to have $300,000 in 20 years.

In case of any query, kindly comment on the solution.